# Unraveling the Smoothness Properties of Diffusion Models: A Gaussian Mixture Perspective

### v1

@@ 5.1 Key definitions and assumptions
Let $p,q$ are two probability distributions, then we have
$\mathrm{TV(p,q)}\leq\sqrt{\frac{1}{2}\mathrm{KL}(p\|q)}.$
@@ 5.2 Total variation
Now, we are ready to present our result for total variation bound assuming data distribution is $k$-mixture of Gaussian ($p_{0}$ in Eq. (9)). Recall $\epsilon_{0}$ is defined in Assumption 5.1, $\epsilon$ is denied in Definition 5.2 and $h_{k}$ is defined in Definition 4.2. See the proof in Appendix I.
Theorem 5.6 ($\mathsf{DDPM}$, total variation, informal version of Theorem I.1).
@@ 5.3 KL divergence
Similarly, we can present our result for the KL divergence bound in the following two theorems, assuming data distribution is $k$-mixture of Gaussian. Recall $\epsilon_{0}$ is defined in Assumption 5.1, $\epsilon$ is denied in Definition 5.2 and $h_{k}$ is defined in Definition 4.2. See the proof in Appendix I.
Theorem 5.7 ($\mathsf{DDPM}$, KL divergence, informal version of Theorem I.2).
@@
where $L=\frac{1}{\sigma_{\min(p_{t})}}+\frac{2R^{2}}{\gamma^{2}\sigma_{\min(p_{t})}^{2}}\cdot(\frac{1}{(2\pi)^{d}\det_{\min(p_{t})}}+\frac{1}{(2\pi)^{d/2}\det_{\min(p_{t})}^{1/2}})\cdot\exp(-\frac{\beta^{2}}{2\sigma_{\max(p_{t})}})$ and $M_{2}=\sum_{i=1}^{k}\alpha_{i}(\|\mu_{i}\|_{2}^{2}+\operatorname{tr}[\Sigma_{i}])$.
Theorem 5.8 ($\mathsf{DDPM}$, KL divergence for smooth data distribution, informal version of Theorem I.3).
@@
where $L=\frac{1}{\sigma_{\min(p_{0})}}+\frac{2R^{2}}{\gamma^{2}(\sigma_{\min(p_{0})})^{2}}\cdot(\frac{1}{(2\pi)^{d}\det_{\min(p_{0})}}+\frac{1}{(2\pi)^{d/2}(\det_{\min(p_{0})})^{1/2}})\cdot\exp(-\frac{\beta^{2}}{2\sigma_{\max(p_{0})}})$ and $M_{2}=\sum_{i=1}^{k}\alpha_{i}(\|\mu_{i}\|_{2}^{2}+\operatorname{tr}[\Sigma_{i}])$. In particular, if we set $T=\Theta(\log(M_{2}d/\epsilon_{0}))$ and $N=\Theta(d^{2}(T+\log L)^{2}/\epsilon_{0}^{2})$, then $\mathrm{KL}(p_{0}\|\widehat{q})\leq\widetilde{O}(\epsilon_{0}^{2})$. In addition, for Euler-Maruyama scheme defined in Definition 4.3, the same bounds hold with an additional $M_{2}\sum_{k=1}^{N}h_{k}^{3}$ term.
@@ 5.4 Probability flow ODE
Theorem 5.9 ($\mathsf{DPOM}$, informal version of Theorem I.4).
@@
where $L=\frac{1}{\sigma_{\min(p_{t})}}+\frac{2R^{2}}{\gamma^{2}\sigma_{\min(p_{t})}^{2}}\cdot(\frac{1}{(2\pi)^{d}\det_{\min(p_{t})}}+\frac{1}{(2\pi)^{d/2}\det_{\min(p_{t})}^{1/2}})\cdot\exp(-\frac{\beta^{2}}{2\sigma_{\max(p_{t})}})$ and $m_{2}=(\sum_{i=1}^{k}\alpha_{i}(\|\mu_{i}\|_{2}^{2}+\operatorname{tr}[\Sigma_{i}]))^{1/2}$. In particular, if we set $T=\Theta(\log(dm_{2}/\epsilon))$, $h_{\mathrm{pred}}=\widetilde{\Theta}(\frac{\epsilon}{L^{2}d^{1/2}})$, $h_{\mathrm{corr}}=\widetilde{\Theta}(\frac{\epsilon}{L^{3}d})$, and if the score estimation error satisfies $\epsilon_{0}\leq\widetilde{O}(\frac{\epsilon}{\sqrt{L}})$, then we can obtain TV error $\epsilon$ with a total iteration complexity of $\widetilde{\Theta}(L^{3}d/\epsilon^{2})$ steps.
Theorem 5.10 ($\mathsf{DPUM}$, informal version of Theorem I.5).
@@ Acknowledgement
  • Section D provides the tools we use form previous papers.
  • Section E provides the case when we consider a continuous time score function, specifically a single Gaussian.
  • Section F provides the case when we consider the score function to be 2 mixtures of Gaussians.
  • Section G provides the case when we consider the score function to be $k$ mixture of Gaussians.
  • Section H provides lemmas that we use for a more concrete calculation for theorems in Section D.
  • Section I provides our main results when we consider the data distribution is $k$ mixture of Gaussians.
@@ C.3 Lipschitz multiplication property
Appendix D Tools From Previous Work
In this section we present several tools we use from previous work.
Assumption D.1 (Lipschitz score, Assumption 1 in [CCL+22], page 6).
For all $t\geq 0$, the score $\nabla\log p_{t}$ is $L$-Lipschitz.
Assumption D.2 (Second momentum bound, Assumption 2 in [CCL+22], page 6 and Assumption 2 in [CLL23], page 6).
Assumption D.3 (Smooth data distributions, Assumption 4 in [CLL23], page 10).
Remark D.4.
We state a tool from previous work [CCL+22],
Theorem D.5 (DDPM, Theorem 2 in [CCL+22], page 7).
Suppose that Assumptions D.1, D.2 and 5.1 hold. Let $\widehat{q}_{T}$ be the output of DDPM algorithm at times $T$, and suppose that the step size $h:=T/N$ satisfies $h\lesssim 1/L$, where $L\geq 1$. Then, it holds that
We state a tool from previous work [CLL23].
Theorem D.6 (Theorem 1 in [CLL23]).
Suppose that Assumptions D.1, D.2, 5.1 hold. If $L\geq 1$, $h_{k}\leq 1$ for $k=1,\dots,N$ and $T\geq 1$, using uniform discretization points yields the following:
  • Using Exponential Integrator scheme (8), we have
    $\mathrm{KL}(p_{0}\|\widehat{q}_{T})\lesssim(M_{2}+d)e^{-T}+T\epsilon_{0}^{2}+\frac{dT^{2}L^{2}}{N}.$
  • Using the Euler-Maruyama scheme (7), we have
    $\mathrm{KL}(p_{0}\|\widehat{q}_{T})\lesssim(M_{2}+d)e^{-T}+T\epsilon_{0}^{2}+\frac{dT^{2}L^{2}}{N}+\frac{T^{3}M_{2}}{N^{2}}.$
Theorem D.7 (Theorem 5 in [CLL23], page 10).
There is a universal constant $K$ such that the following holds. Under Assumptions D.2, 5.1, and D.3 hold, by using the exponentially decreasing (then constant) step size $h_{k}=c\min\{\max\{t_{k},\frac{1}{L}\},1\},c=\frac{T+\log L}{N}\leq\frac{1}{Kd}$, the sampling dynamic (8) results in a distribution $\widehat{q}_{T}$ such that
We state a tool from previous work [CCL+23],
Theorem D.8 (DPOM, Theorem 2 in [CCL+23], page 6).
Suppose that Assumptions D.1, D.2 and 5.1 hold. If $\widehat{q}_{T}$ denotes the output of DPOM (see Algorithm 1 in [CCL+23]) with early stopping. Then, it holds that
Theorem D.9 (DPUM, Theorem 3 in [CCL+23], page 7).
Suppose that Assumptions D.1, D.2 and 5.1 hold. If $\widehat{q}_{T}$ denotes the output of DPUM (see Algorithm 2 in [CCL+23]) with early stopping. Then, it holds that
Appendix E Single Gaussian Case
In this section, we consider the continuous case of $p_{t}(x)$, which is the probability density function ($\mathsf{pdf}$) of the input data $x$, and also a function of time $t$. More specifically, we consider the cases when $p_{t}(x)$ is: (1) a single Gaussian where either the mean is a function of time (Section E.1) or the covariance is a function of time (Section E.2), (2) a single Gaussian where both the mean and the covariance are a function of time (Section E.3). And then, we compute the upper bound and Lipschitz constant for the score function i.e. the gradient of log $\mathsf{pdf}$ $\frac{\mathrm{d}\log p_{t}(x)}{\mathrm{d}x}$.
E.1 Case when the mean of $p_{t}(x)$ is a function of time
@@
where the first step follows from Definition E.1, the second step follows from variables are independent, the third step follows from Fact C.1, and the last step follows from simple algebra. ∎
@@
Lemma E.3 (Linear growth).
@@
where the first step follows from Lemma E.2, the second step follows from Fact C.2, and the last step follows from simple algebra. ∎
@@
where the first step follows from Lemma E.2, and the last step follows from simple algebra. ∎
E.2 Case when the covariance of $p_{t}(x)$ is a function of time
@@
where the first step follows from Definition E.5, the second step follows from variables are independent, the third step follows from Fact C.1, and the last step follows from simple algebra. ∎
@@
Lemma E.7 (Linear growth).
@@
where the first step follows from Lemma E.6, the second step follows from Fact C.2, the third step follows from $t\geq 0$, the fourth step follows from Fact C.2, and the last step follows from simple algebra. ∎
@@
where the first step follows from Lemma E.6, the second step follows from simple algebra, the third step follows from Fact C.2. ∎
E.3 A general version for single Gaussian
@@
Lemma E.11 (Linear growth).
@@
where the first step follows from Lemma E.10, the second step follows from Fact C.2, the third step follows from Fact C.2, the fourth step follows from Fact C.2, the fifth step follows from Fact C.2, and the last step follows from Fact C.2. ∎
@@
where the first step follows from Lemma E.10, the second step follows from simple algebra, the third step follows from Fact C.2, and the last step follows from Fact C.2. ∎
Appendix F A General Version for Two Gaussian
@@
  • Section F.1 defines the probability density function ($\mathsf{pdf}$) $p_{t}(x)$ that we use, which is a mixture of 2 Gaussian.
  • Section F.2 provides lemmas that are used for calculation of the score function i.e. gradient of the log $\mathsf{pdf}$ $\frac{\mathrm{d}\log p_{t}(x)}{\mathrm{d}x}$.
  • Section F.3 provides the expression of the score function.
  • Section F.4 provides lemmas that are used for calculation of the upper bound of the score function.
  • Section F.5 provides the expression of the upper bound of the score function.
  • Section F.6 provides lemmas of upper bound for some base functions that are used for calculation of the Lipschitz constant of the score function.
  • Section F.7 provides lemmas of Lipschitz constant for some base functions that are used for calculation of the Lipschitz constant of the score function.
  • Section F.8 provides lemmas of Lipschitz constant for $f(x)$ that are used for calculation of the Lipschitz constant of the score function.
  • Section F.9 provides lemmas of Lipschitz constant for $g(x)$ that are used for calculation of the Lipschitz constant of the score function.
  • Section F.10 provides the expression of the Lipschitz constant of the score function.

### v2

@@ 5.1 Key definitions and assumptions
In Condition 5.3, we denote $\sigma_{\min(p_{0})},\sigma_{\max(p_{0})}$ as the minimum/maximum singular value between the covariance matrices of $k$-components at the original data distribution $p_{0}$. $\mu_{\max(p_{0})}$ is the maximum $\ell_{2}^{2}$-norm between mean vectors of the $k$-components at $p_{0}$. $\det_{\min(p_{0})}$ is the minimum determinant between the covariance matrices of $k$-components at $p_{0}$.
@@
In Condition 5.4, we denote $\sigma_{\min(p_{t})},\sigma_{\max(p_{t})}$ as the minimum/maximum singular value between the covariance matrices of $k$-components at $p_{t}$. $\beta,R$ be the lower/upper bound of the $\ell_{2}$ distance between the input data $x$ and all $k$ scaled mean vectors $a_{t}\mu_{i}$ at timestep $t$. Let $\gamma$ be the lower bound of $p_{t}(x)$, the $\mathsf{pdf}$ at time $t$ and $\det_{\min(p_{t})}$ be the minimum determinant at $p_{t}$.
@@
Let $p,q$ are two probability distributions, then we have $\mathrm{TV(p,q)}\leq\sqrt{\frac{1}{2}\mathrm{KL}(p\|q)}.$
@@ 5.2 Total variation
Now, we are ready to present our result for total variation bound assuming data distribution is $k$-mixture of Gaussian ($p_{0}$ in Eq. (9)). Recall $\epsilon_{0}$ is defined in Assumption 5.1, $\epsilon$ is denied in Definition 5.2 and $h_{k}$ is defined in Definition 4.2. See the proof in Appendix H.
Theorem 5.6 ($\mathsf{DDPM}$, total variation, informal version of Theorem H.1).
@@ 5.3 KL divergence
Similarly, we can present our result for the KL divergence bound in the following two theorems, assuming data distribution is $k$-mixture of Gaussian. Recall $\epsilon_{0}$ is defined in Assumption 5.1, $\epsilon$ is denied in Definition 5.2 and $h_{k}$ is defined in Definition 4.2. See the proof in Appendix H.
Theorem 5.7 ($\mathsf{DDPM}$, KL divergence, informal version of Theorem H.2).
@@
$\mathrm{KL}(p_{0}\|\widehat{q})\lesssim(M_{2}+d)\exp(-T)+T\epsilon_{0}^{2}+\frac{dT^{2}L^{2}}{N}+\frac{T^{3}M_{2}}{N^{2}}.$
where $M_{2}=\sum_{i=1}^{k}\alpha_{i}(\|\mu_{i}\|_{2}^{2}+\operatorname{tr}[\Sigma_{i}])$ and $L=\frac{1}{\sigma_{\min(p_{t})}}+\frac{2R^{2}}{\gamma^{2}\sigma_{\min(p_{t})}^{2}}\cdot(\frac{1}{(2\pi)^{d}\det_{\min(p_{t})}}+\frac{1}{(2\pi)^{d/2}\det_{\min(p_{t})}^{1/2}})\cdot\exp(-\frac{\beta^{2}}{2\sigma_{\max(p_{t})}})$.
Theorem 5.8 ($\mathsf{DDPM}$, KL divergence for smooth data distribution, informal version of Theorem H.3).
@@
where $M_{2}=\sum_{i=1}^{k}\alpha_{i}(\|\mu_{i}\|_{2}^{2}+\operatorname{tr}[\Sigma_{i}])$ and $L=\frac{1}{\sigma_{\min(p_{t})}}+\frac{2R^{2}}{\gamma^{2}\sigma_{\min(p_{t})}^{2}}\cdot(\frac{1}{(2\pi)^{d}\det_{\min(p_{t})}}+\frac{1}{(2\pi)^{d/2}\det_{\min(p_{t})}^{1/2}})\cdot\exp(-\frac{\beta^{2}}{2\sigma_{\max(p_{t})}})$. In particular, if we set $T=\Theta(\log(M_{2}d/\epsilon_{0}))$ and $N=\Theta(d^{2}(T+\log L)^{2}/\epsilon_{0}^{2})$, then $\mathrm{KL}(p_{0}\|\widehat{q})\leq\widetilde{O}(\epsilon_{0}^{2})$. In addition, for Euler-Maruyama scheme defined in Definition 4.3, the same bounds hold with an additional $M_{2}\sum_{k=1}^{N}h_{k}^{3}$ term.
@@ 5.4 Probability flow ODE
Theorem 5.9 ($\mathsf{DPOM}$, informal version of Theorem H.4).
@@
where $m_{2}=(\sum_{i=1}^{k}\alpha_{i}(\|\mu_{i}\|_{2}^{2}+\operatorname{tr}[\Sigma_{i}]))^{1/2}$ and $L=\frac{1}{\sigma_{\min(p_{t})}}+\frac{2R^{2}}{\gamma^{2}\sigma_{\min(p_{t})}^{2}}\cdot(\frac{1}{(2\pi)^{d}\det_{\min(p_{t})}}+\frac{1}{(2\pi)^{d/2}\det_{\min(p_{t})}^{1/2}})\cdot\exp(-\frac{\beta^{2}}{2\sigma_{\max(p_{t})}})$. In particular, if we set $T=\Theta(\log(dm_{2}/\epsilon))$, $h_{\mathrm{pred}}=\widetilde{\Theta}(\frac{\epsilon}{L^{2}d^{1/2}})$, $h_{\mathrm{corr}}=\widetilde{\Theta}(\frac{\epsilon}{L^{3}d})$, and if the score estimation error satisfies $\epsilon_{0}\leq\widetilde{O}(\frac{\epsilon}{\sqrt{L}})$, then we can obtain TV error $\epsilon$ with a total iteration complexity of $\widetilde{\Theta}(L^{3}d/\epsilon^{2})$ steps.
Theorem 5.10 ($\mathsf{DPUM}$, informal version of Theorem H.5).
@@
6 Tools From Previous Work
In this section we present several tools we use from previous work.
Assumption 6.1 (Lipschitz score, Assumption 1 in [CCL+22], page 6).
For all $t\geq 0$, the score $\nabla\log p_{t}$ is $L$-Lipschitz.
Assumption 6.2 (Second momentum bound, Assumption 2 in [CCL+22], page 6 and Assumption 2 in [CLL23], page 6).
Assumption 6.3 (Smooth data distributions, Assumption 4 in [CLL23], page 10).
Remark 6.4.
We state a tool from previous work [CCL+22],
Theorem 6.5 (DDPM, Theorem 2 in [CCL+22], page 7).
Suppose that Assumptions 6.1, 6.2 and 5.1 hold. Let $\widehat{q}_{T}$ be the output of DDPM algorithm at times $T$, and suppose that the step size $h:=T/N$ satisfies $h\lesssim 1/L$, where $L\geq 1$. Then, it holds that
We state a tool from previous work [CLL23].
Theorem 6.6 (Theorem 1 in [CLL23]).
Suppose that Assumptions 6.1, 6.2, 5.1 hold. If $L\geq 1$, $h_{k}\leq 1$ for $k=1,\dots,N$ and $T\geq 1$, using uniform discretization points yields the following:
  • Using Exponential Integrator scheme (8), we have
    $\mathrm{KL}(p_{0}\|\widehat{q}_{T})\lesssim(M_{2}+d)\exp(-T)+T\epsilon_{0}^{2}+\frac{dT^{2}L^{2}}{N}.$
  • Using the Euler-Maruyama scheme (7), we have
    $\mathrm{KL}(p_{0}\|\widehat{q}_{T})\lesssim(M_{2}+d)\exp(-T)+T\epsilon_{0}^{2}+\frac{dT^{2}L^{2}}{N}+\frac{T^{3}M_{2}}{N^{2}}.$
Theorem 6.7 (Theorem 5 in [CLL23], page 10).
There is a universal constant $K$ such that the following holds. Under Assumptions 6.2, 5.1, and 6.3 hold, by using the exponentially decreasing (then constant) step size $h_{k}=c\min\{\max\{t_{k},\frac{1}{L}\},1\},c=\frac{T+\log L}{N}\leq\frac{1}{Kd}$, the sampling dynamic (8) results in a distribution $\widehat{q}_{T}$ such that
We state a tool from previous work [CCL+23],
Theorem 6.8 (DPOM, Theorem 2 in [CCL+23], page 6).
Suppose that Assumptions 6.1, 6.2 and 5.1 hold. If $\widehat{q}_{T}$ denotes the output of DPOM (see Algorithm 1 in [CCL+23]) with early stopping. Then, it holds that
Theorem 6.9 (DPUM, Theorem 3 in [CCL+23], page 7).
Suppose that Assumptions 6.1, 6.2 and 5.1 hold. If $\widehat{q}_{T}$ denotes the output of DPUM (see Algorithm 2 in [CCL+23]) with early stopping. Then, it holds that
@@ Acknowledgement
  • Section D provides the case when we consider a continuous time score function, specifically a single Gaussian.
  • Section E provides the case when we consider the score function to be 2 mixtures of Gaussians.
  • Section F provides the case when we consider the score function to be $k$ mixture of Gaussians.
  • Section G provides lemmas that we use for a more concrete calculation for theorems in Section 6.
  • Section H provides our main results when we consider the data distribution is $k$ mixture of Gaussians.
@@ C.3 Lipschitz multiplication property
Appendix D Single Gaussian Case
In this section, we consider the continuous case of $p_{t}(x)$, which is the probability density function ($\mathsf{pdf}$) of the input data $x$, and also a function of time $t$. More specifically, we consider the cases when $p_{t}(x)$ is: (1) a single Gaussian where either the mean is a function of time (Section D.1) or the covariance is a function of time (Section D.2), (2) a single Gaussian where both the mean and the covariance are a function of time (Section D.3). And then, we compute the upper bound and Lipschitz constant for the score function i.e. the gradient of log $\mathsf{pdf}$ $\frac{\mathrm{d}\log p_{t}(x)}{\mathrm{d}x}$.
D.1 Case when the mean of $p_{t}(x)$ is a function of time
@@
where the first step follows from Definition D.1, the second step follows from variables are independent, the third step follows from Fact C.1, and the last step follows from simple algebra. ∎
@@
Lemma D.3 (Linear growth).
@@
where the first step follows from Lemma D.2, the second step follows from Fact C.2, and the last step follows from simple algebra. ∎
@@
where the first step follows from Lemma D.2, and the last step follows from simple algebra. ∎
D.2 Case when the covariance of $p_{t}(x)$ is a function of time
@@
where the first step follows from Definition D.5, the second step follows from variables are independent, the third step follows from Fact C.1, and the last step follows from simple algebra. ∎
@@
Lemma D.7 (Linear growth).
@@
where the first step follows from Lemma D.6, the second step follows from Fact C.2, the third step follows from $t\geq 0$, the fourth step follows from Fact C.2, and the last step follows from simple algebra. ∎
@@
where the first step follows from Lemma D.6, the second step follows from simple algebra, the third step follows from Fact C.2. ∎
D.3 A general version for single Gaussian
@@
where the first step follows from Definition D.9, the second step follows from Fact C.1, the third step follows from Fact C.3, and the last step follows from simple algebra. ∎
@@
Lemma D.11 (Linear growth).
@@
where the first step follows from Lemma D.10, the second step follows from Fact C.2, the third step follows from Fact C.2, the fourth step follows from Fact C.2, the fifth step follows from Fact C.2, and the last step follows from Fact C.2. ∎
@@
where the first step follows from Lemma D.10, the second step follows from simple algebra, the third step follows from Fact C.2, and the last step follows from Fact C.2. ∎
Appendix E A General Version for Two Gaussian
@@
  • Section E.1 defines the probability density function ($\mathsf{pdf}$) $p_{t}(x)$ that we use, which is a mixture of 2 Gaussian.
  • Section E.2 provides lemmas that are used for calculation of the score function i.e. gradient of the log $\mathsf{pdf}$ $\frac{\mathrm{d}\log p_{t}(x)}{\mathrm{d}x}$.
  • Section E.3 provides the expression of the score function.
  • Section E.4 provides lemmas that are used for calculation of the upper bound of the score function.
  • Section E.5 provides the expression of the upper bound of the score function.
  • Section E.6 provides lemmas of upper bound for some base functions that are used for calculation of the Lipschitz constant of the score function.
  • Section E.7 provides lemmas of Lipschitz constant for some base functions that are used for calculation of the Lipschitz constant of the score function.
  • Section E.8 provides lemmas of Lipschitz constant for $f(x)$ that are used for calculation of the Lipschitz constant of the score function.
  • Section E.9 provides lemmas of Lipschitz constant for $g(x)$ that are used for calculation of the Lipschitz constant of the score function.
  • Section E.10 provides the expression of the Lipschitz constant of the score function.